OpenStudy (anonymous):

At the outer edge of a rotating space habitat, 130 m from its center, the rotational acceleration is g. What is the rotational acceleration at a distance of 65 m from the center of the habitat?

OpenStudy (anonymous):

Is "g" one G-force? Or do you need to give the answer in terms of g?

OpenStudy (anonymous):

(do you need to use 'g' as a variable)?

OpenStudy (anonymous):

am not shure

OpenStudy (anonymous):

I'm sorry, I just realised that it makes no difference. Have you seen the formula: \[a_c = \frac{v^2}{r}\]

OpenStudy (anonymous):

first time seeing it

OpenStudy (anonymous):

Okay, to me this is a physics question but there may well be a more mathematical way of doing it. I will show you the way that I know

OpenStudy (anonymous):

its a physics question

OpenStudy (anonymous):

\[a_c=\frac{v^2}{r}\] From the first part of the statement, we know the acceleration and radius so we can fill that in for the formula. \[g=\frac{v^2}{130}\] Therefore \[v^2=130g\] Now we know the value of v^2, and from the second part of the statement we know the radius and that means we can use this to solve for the acceleration. \[a_c=\frac{v^2}{r}\] \[a_c=\frac{130g}{65}\] \[a_c=2g\] Therefore the acceleration is 2g, or 2Gs (if g is a unit)

OpenStudy (anonymous):

For specifically physics questions, there is a physics part of open study that may be able to help you more. I have just taken physics in highschool so I knew how to do it :)

OpenStudy (anonymous):

thank you :) could you help me out with one more question?

OpenStudy (anonymous):

I'll do my best

OpenStudy (anonymous):

I have to go now, sorry. Asking in the physics openstudy section will probably give you the best luck

OpenStudy (anonymous):

Sorry computer kept on freezing